The selling price of a car is $21,000. If the markup is 20% of the dealer’s cost, what is the dealer’s cost of the car?
Answer by robertb(5830)

21000 = C + 0.20C = 1.20C
==> 21000 = 1.20C
==> C = 21000/1.20 = $17,500